OpenStudy (anonymous):

I have to finish quick. Need help asap. A $5000 investment earns $1575 simple interest in three years. What is the annual interest rate? A. 1.05% B. 3.15% C. 10.5% D. 31.5%

OpenStudy (anonymous):

500 * 3 = 15000 1575 = 15000r 1575 / 15000 = 0.105 0.105 * 100 = 10.5% r = 10.5%

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

No problem